# Seeding neural network quantum states with tensor network states

### v1

@@ Abstract
We find an efficient approach to approximately convert matrix product states (MPSs) into restricted Boltzmann machine wave functions consisting of a multinomial hidden unit through a canonical polyadic (CP) decomposition of the MPSs. This method allows us to generate well-behaved initial neural network quantum states for quantum many-body ground-state calculations in polynomial time of the number of variational parameters and systematically shorten the distance between the initial states and the ground states with increasing the rank of the CP decomposition. We demonstrate the efficiency of our method by taking the transverse-field Ising model as an example and discuss possible applications of our method to more general quantum many-body systems in which the ground-state wave functions possess complex nodal structures.
@@ I Introduction
Recently, neural network quantum states (NNQSs) have been proposed as a new class of variational wave functions for quantum many-body systems [19, 20, 21, 22, 23, 24, 25, 26]. One of the basic NNQSs is the restricted Boltzmann machine (RBM) wave function [19, 27, 28, 29, 30, 31, 32, 33, 34]. In contrast to tensor network states, NNQSs try to represent quantum many-body states by overparameterizing variational wave functions with neural networks. Because of the flexibility of the network structure and the large number of variational parameters, NNQSs can, in principle, represent arbitrary quantum many-body states even with the volume-law entanglement entropy [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], irrespective of the spatial dimensionality of quantum systems [47, 48, 49].
@@ II.2 CP decomposition of tensors
where $R$ is the rank of the CP decomposition and $A^{(j)}_{i_{j},r}\in\mathbb{C}^{d_{i}\times R}$ is the $r$th component of the vector corresponding to the $j$th axis. The CP decomposition becomes more accurate as the rank $R$ increases. We will discuss the case where the tensor is represented as MPSs in the next section.
@@
$A_{i,r}$ $=\sum_{j,k}X_{i,j,k}\sum_{r^{\prime}}B^{*}_{j,r^{\prime}}C^{*}_{k,r^{\prime}}\left(\sum_{j^{\prime},k^{\prime}}B^{*}_{j^{\prime},r^{\prime}}B_{j^{\prime},r}C^{*}_{k^{\prime},r^{\prime}}C_{k^{\prime},r}\right)^{-1}$ (5)
$=X_{(1)}\left(C^{*}\odot B^{*}\right)\left(C^{\dagger}C\circledast B^{\dagger}B\right)^{-1},$ (6)
@@
$=X_{(2)}\left(C^{*}\odot A^{*}\right)\left(C^{\dagger}C\circledast A^{\dagger}A\right)^{-1}.$ (8)
@@
$C_{k,r}$ $=\sum_{i,j}X_{i,j,k}\sum_{r^{\prime}}A^{*}_{i,r^{\prime}}B^{*}_{j,r^{\prime}}\left(\sum_{i^{\prime},j^{\prime}}A^{*}_{i^{\prime},r^{\prime}}A_{i^{\prime},r}B^{*}_{j^{\prime},r^{\prime}}B_{j^{\prime},r}\right)^{-1}$ (9)
$=X_{(3)}\left(A^{*}\odot B^{*}\right)\left(A^{\dagger}A\circledast B^{\dagger}B\right)^{-1}.$ (10)
@@
$A^{(j)}_{i_{j},r}$ $=\sum_{r^{\prime}}\frac{\sum_{i_{1},i_{2},\dots,i_{j-1},i_{j+1},\dots,i_{n}}X_{i_{1},i_{2},\dots,i_{n}}\prod_{k=1,k\neq j}^{n}A^{(k)*}_{i_{k},r^{\prime}}}{\prod_{k^{\prime}=1,k^{\prime}\neq j}^{n}\sum_{i^{\prime}_{k^{\prime}}}A^{(k^{\prime})*}_{i^{\prime}_{k^{\prime}},r^{\prime}}A^{(k^{\prime})}_{i^{\prime}_{k^{\prime}},r}}.$ (12)
@@ II.3 CP decomposition of MPSs
$\delta_{r_{1},r_{2},\dots,r_{n}}$ =$=\begin{cases}1&(r_{1}=r_{2}=\cdots=r_{n}),\\ 0&(\text{otherwise}),\end{cases}$ (13)
@@
$X_{i_{1},i_{2},\dots,i_{n}}$ =$=\sum_{\{\alpha_{i}\}}F^{(1)}_{\alpha_{0},i_{1},\alpha_{1}}F^{(2)}_{\alpha_{1},i_{2},\alpha_{2}}\cdots F^{(n)}_{\alpha_{n-1},i_{n},\alpha_{n}},$ (15)
@@
The counterpart of the $d_{i}\times R$ matrix $X_{(1)}(C^{*}\odot B^{*})$, as shown in Fig. 3(b), can be evaluated in the following manner. For factors $A^{(j)}$ ($12\leq j\leq n-$ 1) having the MPS representations $F^{(i)}$, we construct the left and right matrices $L^{(j)}$ and $R^{(j)}$, defined as
@@
$~{}\cdot G^{(2)}_{\alpha_{1},s_{2},\alpha_{2}}F^{(2)*}_{\tilde{\alpha}_{1},s_{2},\tilde{\alpha}_{2}}\cdots G^{(j-1)}_{\alpha_{j-2},s_{j-1},\alpha_{j-1}},F^{(j-1)*}_{\tilde{\alpha}_{j-2},s_{j-1},\tilde{\alpha}_{j-1}},$ (17)
@@
respectively, as shown in Figs. 3(c,d). Then, we can evaluate the $d_{i}\times R$ matrix $Y^{(j)}$ as
@@
For factor $A^{(1)}$ [$A^{(n)}$], we only need the right (left) matrix to construct matrix $Y^{(j)}$.
The counterpart of the $R\times R$ matrix $(C^{\dagger}C\circledast B^{\dagger}B)$ can be evaluated just as in the original ALS method. We have to calculate the element-wise product of the $R\times R$ matrices $n-2$ times. The corresponding matrix $Z^{(j)}$ is given by
@@ II.4 RBM wave functions with multinomial hidden units derived from the CP decomposed MPSs
Here, $a_{i}$, $b_{j}$, and $W_{i,j}$ are the parameters of the RBM. The symbols $n_{v}$ and $n_{h}$ represent the number of the visible and hidden units, respectively. The variable $v_{i}(=s_{i})$ is the index of the $i$th spin and takes the value $+1$ or $-1$, depending on the state of the spin $\uparrow$ or $\downarrow$. For binomial hidden units, the variable $h_{j}$ also takes the value $+1$ or $-1$. Note that, for simplicity, we ignore the normalization factor of the wave function amplitude, which corresponds to the partition function $Z=\sum_{\{v_{i}\},\{h_{j}\}}\exp[-E(\{v_{i}\},\{h_{j}\})]$.
@@
where only the $k$th component is $1$ and all other components are 00. Then, the virtual energy function is given by
@@
$\Psi(\{v_{i}\})$ =$=\sum_{\{h_{j}\}}\exp\left(\sum_{i=1}^{n_{v}}a_{i}v_{i}+\sum_{j=1}^{n_{h}}\sum_{k=1}^{n_{r}}h_{j,k}\theta_{j,k}\right)$ (28)
@@
$\frac{b_{j=1,k}}{n_{v}}$ $=\frac{\mathrm{Log\,}A^{(l)}_{i_{l}=0,k}+\mathrm{Log\,}A^{(l)}_{i_{l}=1,k}}{2},$ (37)
@@
In general, as the number of hidden units increases, the RBM wave function well approximates the ground state of the quantum many-body system, which is known as the universal approximation theorem [88]. By contrast, as for the present RBM wave function with a single multinomial hidden unit, this theorem does not necessarily hold. However, by definition, the CP decomposition exactly reproduces the original tensor when $R$ is larger than or equal to the CP rank of the tensor [73]; even for $R$ that is smaller than the CP rank, the CP decomposition would nearly reconstruct the original tensor when $R$ is sufficiently large [74, 75]. This fact implies that the RBM wave function with a single hidden unit has a potential to represent the ground state of the quantum many-body system when $R$ is sufficiently large, although $R$ could be exponentially large in the number of sites $n$ in the worst case.
@@ II.5 Optimization of RBM wave functions
The right-hand side of the above equations is expressed by the softmax function. To avoid numerical instabilities such as overflow and underflow, we typically calculate a maximum of a real part of $\tilde{\theta}_{k}$ [$m=\max_{k}\mathrm{Re\,}\tilde{\theta}_{k}$] and estimate the exponentials in the numerator and denominator by $\exp(\tilde{\theta}_{k}-m)$. The most time-consuming part of the SR method is solving the linear equations with a large number of variational parameters. In general, the computational cost scales quadratically with the number of variational parameters [19].
@@ III Applications to quantum many-body problems: transverse-field Ising model
The ground state of the one-dimensional transverse-field Ising model can be solved analytically using the Jordan-Wigner transformation [95]. The ground-state energy is given by
@@
For periodic boundary conditions with even total fermion number, the summation over $k$ is taken over the set of momenta, satisfying
@@
Hereafter, we first consider the model under open boundary conditions and obtain the ground-state wave function in the MPS representation using the DMRG method. We apply the TeNPy library for the DMRG method [6, 7] and choose the bond dimension that is large enough to represent the true ground-state energy within a desired accuracy, which is smaller than $10^{-9}$ in the unit of $J$ in this study. We turn off the parity conservation during the DMRG simulation to purposely keep the wave function primitive. We then apply the CP decomposition to the MPS representation of the ground state. We find that the CP decomposed MPSs are sufficiently close to the ground-state wave function when the rank of the CP decomposition is large. The infidelity between the ground-state MPS and the approximated CP decomposed tensor decreases at most polynomially as the rank of the CP decomposition increases. We then approximately convert the MPS representation into the RBM wave function consisting of a single hidden unit and apply the VMC method to optimize the parameters in the RBM wave function. We demonstrate that the initial state prepared under open boundary conditions can be efficiently used for the simulation under periodic boundary conditions.
Note that the ground state of the present transverse-field Ising model does not have a complex nodal structure. The amplitude of the ground-state wave function can be taken as nonnegative according to the Perron-Frobenius theorem [29, 98]. This fact suggests that one can use positive factors for the CP decomposition in this study. However, to analyze the performance of the CP decomposition more generally, we purposely apply the CP decomposition allowing both positive and negative factors. As we will see later, the CP decomposition with positive and negative factors also well approximates the ground-state wave function and provides a good initial state having complex amplitudes for the VMC simulation.
@@ III.1 Accuracy of the CP decomposition of MPSs
We calculate two types of errors to evaluate the accuracy of the CP decomposition of MPSs. One is the infidelity ($\bar{F}$) between the ground-state MPS ($|\mathrm{MPS}\rangle$) and the approximated CP decomposed tensor ($|\mathrm{CP}\rangle$), defined by
@@
When the CP decomposition exactly reproduces the ground-state wave function, the infidelity becomes zero. For simplicity, all the elements in $|\mathrm{MPS}\rangle$ and $|\mathrm{CP}\rangle$ are assumed to be real in the transverse-field Ising model. The other is the energy difference between the exact ground-state energy ($E_{\mathrm{exact}}$) and the energy calculated from the approximated CP decomposed tensor ($E_{\mathrm{CP}}$), defined by
@@
We will see that both of these errors decrease polynomially as the rank of the CP decomposition increases.
@@
Figure 5: (a) Infidelity and energy difference as a function of the rank $R$ of the CP decomposition for $h/J=2$ and $n=16$. Different symbols represent different initial factors for the CP decomposition. Open symbols are for the infidelity and filled symbols are for the energy difference. (b) Same as (a) but they are shown in the logarithmic scale. We plot the line proportional to $R^{-2}$ as a reference.
@@
Figure 7: (a) Infidelity and energy difference as a function of the rank $R$ of the CP decomposition for $h/J=0.5$ and $n=16$. Different symbols represent different initial factors for the CP decomposition. Open symbols are for the infidelity and filled symbols are for the energy difference. (b) Same as (a) but they are shown in the logarithmic scale. We plot the line proportional to $R^{-2}$ as a reference.
@@
Figure 9: (a) Infidelity and energy difference as a function of the rank $R$ of the CP decomposition for $h/J=1$ and $n=16$. Different symbols represent different initial factors for the CP decomposition. Open symbols are for the infidelity and filled symbols are for the energy difference. (b) Same as (a) but they are shown in the logarithmic scale. We plot the line proportional to $R^{-2}$ as a reference.
@@
We illustrate the $R$ dependence of the errors in Fig. 5. Both infidelity and energy difference decrease monotonically as $R$ increases. The decay appears to be faster than $\sim R^{-1}$. Indeed, for larger $R$, $R$ dependences of both errors are well approximated by the line proportional to $R^{-2}$.
@@
We also show the $R$ dependence of the errors in Fig. 7. Both infidelity and energy difference decrease rapidly as $R$ increases. When $R\geq 24$, the energy error becomes less than $10^{-4}$ and is almost invisible from the plot. When plotting the errors in the logarithmic scale, we confirm that the decay appears to obey the same $R$ dependence ($\sim R^{-2}$) as the case for $h/J=2$.
@@
We show the $R$ dependence of the errors at the critical point in Fig. 9. Both infidelity and energy difference decrease monotonically as $R$ increases. For larger $R$, $R$ dependences of both errors satisfy the power-law decay $\sim R^{-2}$. The exponent of the decay is consistent with what we observed in the cases for $h/J=0.5$ and $h/J=2$. Both infidelity and energy difference behave in the same way for all parameters $h/J$ that we have considered. Therefore, both quantities can be viewed as an equivalent indicator of the errors.

### v3

@@ Abstract
We find an efficient approach to approximately convert matrix product states (MPSs) into restricted Boltzmann machine wave functions consisting of a multinomial hidden unit through a canonical polyadic (CP) decomposition of the MPSs. This method allows us to generate well-behaved initial neural network quantum states for quantum many-body ground-state calculations in polynomial time of the number of variational parameters and systematically shorten the distance between the initial states and the ground states while increasing the rank of the CP decomposition. We demonstrate the efficiency of our method by taking the transverse-field Ising model as an example and discuss possible applications of our method to more general quantum many-body systems in which the ground-state wave functions possess complex nodal structures.
@@ I Introduction
Recently, neural network quantum states (NNQSs) have been proposed as a new class of variational wave functions for quantum many-body systems [19, 20, 21, 22, 23, 24, 25, 26]. One of the basic NNQSs is the restricted Boltzmann machine (RBM) wave function [19, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. In contrast to tensor network states, NNQSs try to represent quantum many-body states by overparameterizing variational wave functions with neural networks. Because of the flexibility of the network structure and the large number of variational parameters, NNQSs can, in principle, represent arbitrary quantum many-body states even with the volume-law entanglement entropy [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50], irrespective of the spatial dimensionality of quantum systems [51, 52, 53].
@@ II.2 CP decomposition of tensors
where $R$ is the rank of the CP decomposition and $A^{(j)}_{i_{j},r}\in\mathbb{C}^{d_{i}\times R}$ is the $r$th component of the vector corresponding to the $j$th axis. The CP decomposition becomes more accurate as the rank $R$ increases. We discuss the case where the tensor is represented as MPSs in the next section.
@@
$A_{i,r}=\sum_{j,k}X_{i,j,k}\sum_{r^{\prime}}B^{*}_{j,r^{\prime}}C^{*}_{k,r^{\prime}}\left(\sum_{j^{\prime},k^{\prime}}B^{*}_{j^{\prime},r^{\prime}}B_{j^{\prime},r}C^{*}_{k^{\prime},r^{\prime}}C_{k^{\prime},r}\right)^{-1}.$ (5)
The corresponding matrix form is
$A=X_{(1)}\left(C^{*}\odot B^{*}\right)\left(C^{\dagger}C\circledast B^{\dagger}B\right)^{-1},$ (6)
@@
and the matrix form as
$B$ $=X_{(2)}\left(C^{*}\odot A^{*}\right)\left(C^{\dagger}C\circledast A^{\dagger}A\right)^{-1}.$ (8)
@@
$C_{k,r}$ $=\sum_{i,j}X_{i,j,k}\sum_{r^{\prime}}A^{*}_{i,r^{\prime}}B^{*}_{j,r^{\prime}}\left(\sum_{i^{\prime},j^{\prime}}A^{*}_{i^{\prime},r^{\prime}}A_{i^{\prime},r}B^{*}_{j^{\prime},r^{\prime}}B_{j^{\prime},r}\right)^{-1},$ (9)
and the matrix form is given by
$C$ $=X_{(3)}\left(B^{*}\odot A^{*}\right)\left(B^{\dagger}B\circledast A^{\dagger}A\right)^{-1}.$ (10)
@@
$A^{(j)}$ $=X_{(j)}\left[A^{(n)}\odot\cdots\odot A^{(j+1)}\odot A^{(j-1)}\odot\cdots\odot A^{(1)}\right]$
$~\phantom{=}\cdot\bigl[{A^{(n)}}^{\dagger}A^{(n)}\circledast\cdots\circledast{A^{(j+1)}}^{\dagger}A^{(j+1)}$
$~\phantom{=}\circledast{A^{(j-1)}}^{\dagger}A^{(j-1)}\circledast\cdots\circledast{A^{(1)}}^{\dagger}A^{(1)}\bigr]^{-1}.$ (12)
Here, the centered dot $\cdot$ is the conventional matrix product.
@@ II.3 CP decomposition of MPSs
$\delta_{r_{1},r_{2},\dots,r_{n}}$ $=\begin{cases}1&(r_{1}=r_{2}=\cdots=r_{n})\\ 0&(\text{otherwise}),\end{cases}$ (13)
@@
$X_{i_{1},i_{2},\dots,i_{n}}$ $=\sum_{\{\alpha_{i}\}}F^{(1)}_{\alpha_{0},i_{1},\alpha_{1}}F^{(2)}_{\alpha_{1},i_{2},\alpha_{2}}\cdots F^{(n)}_{\alpha_{n-1},i_{n},\alpha_{n}}$, (15)
@@
The counterpart of the $d_{i}\times R$ matrix $X_{(1)}(C^{*}\odot B^{*})$, as shown in Fig. 3(b), can be evaluated in the following manner: For factors $A^{(j)}$ ($2\leq j\leq n-1$) having the MPS representations $F^{(i)}$, we construct the left and right matrices $L^{(j)}$ and $R^{(j)}$, defined as
@@
$~\cdot G^{(2)}_{\alpha_{1},s_{2},\alpha_{2}}F^{(2)*}_{\tilde{\alpha}_{1},s_{2},\tilde{\alpha}_{2}}\cdots G^{(j-1)}_{\alpha_{j-2},s_{j-1},\alpha_{j-1}}F^{(j-1)*}_{\tilde{\alpha}_{j-2},s_{j-1},\tilde{\alpha}_{j-1}}$, (17)
@@
respectively, as shown in Figs. 3(c) and 3(d). Then, we can evaluate the $d_{i}\times R$ matrix $Y^{(j)}$ as
@@
For factor $A^{(1)}$ $A^{(n)}$], we only need the right (left) matrix to construct matrix $Y^{(j)}$.
The counterpart of the $R\times R$ matrix $(C^{\dagger}C\circledast B^{\dagger}B)$ can be evaluated just as in the original ALS method. We have to calculate the element-wise product of the $R\times R$ matrices n2n-2 times. The corresponding matrix $Z^{(j)}$ is given by
@@ II.4 RBM wave functions with multinomial hidden units derived from the CP decomposed MPSs
Here, $a_{i}$, $b_{j}$, and $W_{i,j}$ are the parameters of the RBM. The symbols $n_{v}$ and $n_{h}$ represent the number of the visible and hidden units, respectively. The variable $v_{i}(=s_{i})$ is the index of the $i$th spin and takes the value +1+1 or 1-1, depending on the state of the spin $\uparrow$ or $\downarrow$. For binomial hidden units, the variable $h_{j}$ also takes the value +1+1 or 1-1. Note that, for simplicity, we ignore the normalization factor of the wave function amplitude, which is given by $\sqrt{\langle\Psi|\Psi\rangle}=\sqrt{\sum_{\{s_{i}\}}|\Psi(\{s_{i}\})|^{2}}$.
@@
where only the $k$th component is $1$ and all other components are 0. Then, the virtual energy function is given by
@@
$\Psi(\{v_{i}\})$ $=\sum_{\{h_{j}\}}\exp\left(\sum_{i=1}^{n_{v}}a_{i}v_{i}+\sum_{j=1}^{n_{h}}\sum_{k=1}^{n_{r}}h_{j,k}\theta_{j,k}\right)$ (28)
@@
$\sum_{l=1}^{n_{v}}\frac{b_{j=1,k}}{n_{v}}$ $=\sum_{l=1}^{n_{v}}\frac{\mathrm{Log\,}A^{(l)}_{i_{l}=0,k}+\mathrm{Log\,}A^{(l)}_{i_{l}=1,k}}{2},$ (37)
@@
In general, as the number of hidden units increases, the RBM wave function well approximates the ground state of the quantum many-body system, which is known as the universal approximation theorem [92]. By contrast, as for the present RBM wave function with a single multinomial hidden unit, this theorem does not necessarily hold. However, by definition, the CP decomposition exactly reproduces the original tensor when $R$ is larger than or equal to the CP rank of the tensor [77]; even for $R$ that is smaller than the CP rank, the CP decomposition would nearly reconstruct the original tensor when $R$ is sufficiently large [78, 79]. This fact implies that the RBM wave function with a single hidden unit has a potential to represent the ground state of the quantum many-body system when $R$ is sufficiently large, although $R$ could be exponentially large in the number of sites $n$ in the worst case [93].
@@ II.5 Optimization of RBM wave functions
The right-hand side of the above equations is expressed by the softmax function. To avoid numerical instabilities such as overflow and underflow, we typically calculate a maximum of a real part of $\tilde{\theta}_{k}$ $m=\max_{k}\mathrm{Re\,}\tilde{\theta}_{k}$] and estimate the exponentials in the numerator and denominator by $\exp(\tilde{\theta}_{k}-m)$. The most time-consuming part of the SR method is solving the linear equations with a large number of variational parameters. In general, the computational cost scales quadratically with the number of variational parameters [19].
@@ III Applications to quantum many-body problems: transverse-field Ising model
The ground state of the one-dimensional transverse-field Ising model can be solved analytically using the Jordan-Wigner transformation [100], which maps spin operators onto fermionic creation and annihilation operators. The ground-state energy is given by
@@
Hereafter, the energy is expressed in the units of $J$. For periodic boundary conditions with even total fermion number, the summation over $k$ is taken over the set of momenta, satisfying
@@
Hereafter, we first consider the model under open boundary conditions and obtain the ground-state wave function in the MPS representation using the DMRG method. We apply the TeNPy library for the DMRG method [6, 7] and choose the bond dimension that is large enough to represent the true ground-state energy within a desired accuracy, which is smaller than $10^{-9}$ in the units of $J$ in this study. We turn off the parity conservation during the DMRG simulation to purposely keep the wave function primitive. We then apply the CP decomposition to the MPS representation of the ground state. We find that the CP decomposed MPSs are sufficiently close to the ground-state wave function when the rank of the CP decomposition is large. The infidelity between the ground-state MPS and the approximated CP decomposed tensor decreases at most polynomially as the rank of the CP decomposition increases. We then approximately convert the MPS representation into the RBM wave function consisting of a single hidden unit and apply the VMC method to optimize the parameters in the RBM wave function. We demonstrate that the initial state prepared under open boundary conditions can be efficiently used for the simulation under periodic boundary conditions.
Note that the ground state of the present transverse-field Ising model does not have a complex nodal structure. The amplitude of the ground-state wave function can be taken as non-negative according to the Perron-Frobenius theorem [29, 103]. This fact suggests that one can use positive factors for the CP decomposition in this study. However, to analyze the performance of the CP decomposition more generally, we purposely apply the CP decomposition allowing both positive and negative factors. As we see later, the CP decomposition with positive and negative factors also well approximates the ground-state wave function and provides a good initial state having complex amplitudes for the VMC simulation.
@@ III.1 Accuracy of the CP decomposition of MPSs
We calculate two types of errors to evaluate the accuracy of the CP decomposition of MPSs. One is the infidelity $\bar{F}$ between the ground-state MPS ($|\mathrm{MPS}\rangle$) and the approximated CP decomposed tensor ($|\mathrm{CP}\rangle$), defined by
@@
When the CP decomposition exactly reproduces the ground-state wave function, the infidelity becomes zero. For simplicity, all the elements in $|\mathrm{MPS}\rangle$ and $|\mathrm{CP}\rangle$ are assumed to be real in the transverse-field Ising model. The other is the energy difference between the exact ground-state energy $E_{\mathrm{exact}}$ and the energy calculated from the approximated CP decomposed tensor ($E_{\mathrm{CP}}$), defined by
@@
We see that both of these errors decrease polynomially as the rank of the CP decomposition increases.
@@
Figure 5: (a) Infidelity and energy difference as a function of the rank $R$ of the CP decomposition for $h/J=2$ and $n=16$. Different symbols represent different initial factors for the CP decomposition. Open symbols are for the infidelity and filled symbols are for the energy difference. (b) Same as panel (a) but shown on a logarithmic scale. We plot the line proportional to $R^{-2}$ as a reference. Note that, hereafter, all energy values are expressed in the units of $J$.
@@
Figure 7: (a) Infidelity and energy difference as a function of the rank $R$ of the CP decomposition for $h/J=0.5$ and $n=16$. Different symbols represent different initial factors for the CP decomposition. Open symbols are for the infidelity and filled symbols are for the energy difference. (b) Same as panel (a) but shown on a logarithmic scale. We plot the line proportional to $R^{-2}$ as a reference.
@@
Figure 9: (a) Infidelity and energy difference as a function of the rank $R$ of the CP decomposition for $h/J=1$ and $n=16$. Different symbols represent different initial factors for the CP decomposition. Open symbols are for the infidelity and filled symbols are for the energy difference. (b) Same as panel (a) but shown on a logarithmic scale. We plot the line proportional to $R^{-2}$ as a reference.
@@
We illustrate the $R$ dependence of the errors in Fig. 5. Both infidelity and energy difference decrease monotonically as $R$ increases. The decay appears to be faster than $\sim R^{-1}$. Indeed, for larger $R$, $R$ dependencies of both errors are well approximated by the line proportional to $R^{-2}$.
@@
We also show the $R$ dependence of the errors in Fig. 7. Both infidelity and energy difference decrease rapidly as $R$ increases. When $R\geq 24$, the energy error becomes less than $10^{-4}$ and is almost invisible from the plot. When plotting the errors in the logarithmic scale, we confirm that the decay appears to obey the same $R$ dependence ($\sim R^{-2}$) as is the case for $h/J=2$.
@@
We show the $R$ dependence of the errors at the critical point in Fig. 9. Both infidelity and energy difference decrease monotonically as $R$ increases. For larger $R$, $R$ dependencies of both errors satisfy the power-law decay $\sim R^{-2}$. The exponent of the decay is consistent with what we observed in the cases for $h/J=0.5$ and $h/J=2$. Both infidelity and energy difference behave in the same way for all parameters $h/J$ that we have considered. Therefore, both quantities can be viewed as an equivalent indicator of the errors.